OpenStudy (anonymous):

Suppose you own a card shop. You buy one line of cards at a rate of 4 cards for $5. You plan to sell the cards at a rate of 3 cards for $5. How many cards must you sell in order to make a profit of $100?

OpenStudy (anonymous):

yo motha

OpenStudy (anonymous):

...

OpenStudy (anonymous):

244 cards

OpenStudy (anonymous):

i need to know how

OpenStudy (anonymous):

5/4=1.24 and 5/3=1.66

OpenStudy (anonymous):

i know that

OpenStudy (anonymous):

1.66-1.25=.41

OpenStudy (anonymous):

i want a better solution

OpenStudy (anonymous):

100/.41=243.9

OpenStudy (anonymous):

round to 244

OpenStudy (anonymous):

If you sold 3 3-packs, you would be even with 3 cards left over which you could sell for $5 and make a $5 profit. So, that would come from selling 12 cards (1 leftover from each of the 4-packs you bought) So, if you sell 12 and clear $5, you sell 240 and clear $100. Answer is 240. It's multiplying by 20.

OpenStudy (anonymous):

tcarroll is wrong

OpenStudy (anonymous):

...

OpenStudy (anonymous):

its not 3packs

OpenStudy (anonymous):

i understand him

OpenStudy (anonymous):

but @tcarroll010 what do you mean by "clear"

OpenStudy (anonymous):

my answer right

OpenStudy (anonymous):

Answer is clearly 240. It comes out to 240 on the head.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Clear is a slang business term for "clear profit". You covered your expenses.

OpenStudy (anonymous):

Further explanation:

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i get it now @tcarroll010 Thanks alot!! :D

OpenStudy (anonymous):

When you buy 3 4-packs, you spend $15. When you sell 3 3-packs, you get $15. so you are back to even. But you have 3 cards left over to sell for profit or "gravy". You get that $5 profit when you sell that 4th 3-pack. Selling 12 cards.

OpenStudy (anonymous):

You're quite welcome!